OpenStudy (themathasker):

HELP!!!!!!!! :C Solve the equation c over 4 - 5 = 4

OpenStudy (anonymous):

\[\frac{c}{4-5}=4 \\ \\ -c=4 \\ \\ c=-4\]

OpenStudy (anonymous):

c = 36

OpenStudy (usukidoll):

that could've been \[\frac{c}{4} - 5 = 4\]

OpenStudy (usukidoll):

since it's just c / 4 ? right ? @TheMathAsker is what I typed the correct question?

OpenStudy (usukidoll):

you multiply 4 throughout the entire equation to have c by itself. then add 20 to both sides.

OpenStudy (usukidoll):

c-20=16 c-20+20=16+20 c=36

OpenStudy (themathasker):

Thank you ~! ^^